OpenStudy (anonymous):

d/dx(lnx))/((e^(x)^1/2)(tan^4(4x))

OpenStudy (rogue):

Is this your problem?\[\frac {d}{dx} \left[ \frac {\ln x}{e^{\sqrt {x}} \tan^4 (4x)} \right]\]

OpenStudy (anonymous):

yeah correct

sam (.sam.):

differentiate? or integrate?

OpenStudy (anonymous):

differentiate

OpenStudy (anonymous):

logarithm differentiation

OpenStudy (rogue):

Its quite an ugly problem, involving quotient rule, product rule, chain rule, etc. I'm tempted to try this with logarithmic differentiation.

OpenStudy (rogue):

oh, lol, we have to use log diff.

OpenStudy (anonymous):

lol yeah :)

OpenStudy (anonymous):

am getting